Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probabilitry

Expert replies
by Md Raihan Uddin » Thu Oct 02, 2014 12:51 pm
Three canons are firing at a target. Their individual probabilities to hit the target are 0.3, 0.4 and 0.5. What is the probability that none of the canons would hit the target?
1)0.06
2)0.12
3)0.21
4)0.29
5)0.94
First Method
= 1-(0.3*0.4*0.5)
=0.94
2nd Method
(1-0.3)*(1-0.4)*(1-0.5)=0.7*0.6*0.5=0.21

Now my question is why 1st method is wrong and why 2nd one is correct. How will I understand which one to use and when?
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Thu Oct 02, 2014 1:18 pm
Md Raihan Uddin wrote: First Method
= 1-(0.3*0.4*0.5)
=0.94

Now my question is why 1st method is wrong
The portion in red represents the probability that ALL 3 cannons hit the target.
Thus:
1 - (0.3 * 0.4 * 0.5) = 0.94 = the probability that NOT ALL 3 CANNONS hit the target.

There are 3 ways that NOT ALL 3 CANNONS may hit the target:
Case 1: No cannon hits the target.
Case 2: Exactly 1 cannon hits the target.
Case 3: exactly 2 cannons hit the target.

Thus:
0.94 = P(no cannon hits the target OR exactly 1 cannon hits the target OR exactly 2 cannons hit the target).
Not what the question is asking.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Brent@GMATPrepNow » Thu Oct 02, 2014 2:52 pm
Md Raihan Uddin wrote:Three canons are firing at a target. Their individual probabilities to hit the target are 0.3, 0.4 and 0.5. What is the probability that none of the canons would hit the target?
A) 0.06
B) 0.12
C) 0.21
D) 0.29
E) 0.94
P(1st MISSES target) = 1 - P(1st HITS target) = 1 - 0.3 = 0.7
P(2nd MISSES target) = 1 - P(2nd HITS target) = 1 - 0.4 = 0.6
P(3rd MISSES target) = 1 - P(3rd HITS target) = 1 - 0.5 = 0.5
----------------------------

P(none hit the target) = P(1st misses target AND 2nd misses target AND 3rd misses target)
= P(1st misses target) x P(2nd misses target) x P(3rd misses target)
= 0.7 x 0.6 x 0.5
[spoiler]= 0.21
= C[/spoiler]

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by [email protected] » Thu Oct 02, 2014 10:46 pm
Hi Md Raihan Uddin,

In these types of multi-part probability questions, the details matter a great deal, so you should force yourself to take plenty of notes. In that way, you'll know exactly what each "math step" will be. My guess is that your math skills are just fine, so the issue is really about how you organize your work.

Here, we're told the individual probabilities for three canons to hit a target (.3, .4 and .5, respectively). We're asked for the probability that NONE of the canons hit the target.

From a note-taking standpoint, here is one way that you could write down the question:

(Prob that 1st misses)(Prob that 2nd misses)(Prob that 3rd misses) = ?

From here, you can just input the number and multiply….

(.7)(.6)(.5) = .21

Final Answer: C

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by pooja181 » Thu Oct 02, 2014 10:56 pm
The portion in red represents the probability that ALL 3 cannons hit the target.
Thus:
1 - (0.3 * 0.4 * 0.5) = 0.94 = the probability that NOT ALL 3 CANNONS hit the target.

There are 3 ways that NOT ALL 3 CANNONS may hit the target:
Case 1: No cannon hits the target.
Case 2: Exactly 1 cannon hits the target.
Case 3: exactly 2 cannons hit the target.

Thus:
0.94 = P(no cannon hits the target OR exactly 1 cannon hits the target OR exactly 2 cannons hit the target).
Not what the question is asking.
My skype id- varun289verma

Visit http//
www.freegmat.in

for free gmat practice test - unlimited and free....


'to beat the gmat - take daily test at freegmat.in than analyze each question '
Join the discussion

by Matt@VeritasPrep » Fri Oct 03, 2014 12:11 pm
It often helps to get all your probabilities in order first.

First Cannon:
Hit = .3
Miss = .7

Second Cannon:
Hit = .4
Miss = .6

Third Cannon:
Hit = .5
Miss = .5

If we want ALL THREE cannons to hit, we'd have .3 * .4 * .5.

If we do 1 - (.3 * .4 * .5), we'd have the chances that NOT ALL THREE cannons hit; in other words, the odds that AT LEAST ONE cannon misses.

The important thing here is that the negation of "all" is not "none", it's "not all"! If I say "not all the cannons hit", some of them could still hit the target; there's just (at least) one that misses. This logic will come into play in critical reasoning too, so it's worth practicing.
Join the discussion